# Revised Mass and Orbit of Eridani b: A 1 Jupiter-Mass Planet on a Near-Circular Orbit

### v1

@@ Abstract
The mature Jovian planet Eridani bĀ orbits one of the closest sun-like stars at a moderate separation of 3.5 AU, presenting one of the best opportunities to image a true analog to a solar system planet. We perform a thorough joint reanalysis and cross-validation of all available archival radial velocity and astrometry data, combining data from eight radial velocity instruments and four astrometric sources (Hipparcos, Hubble FGS, Gaia DR2, and Gaia DR3). We incorporate methodological advances that impact our findings including a principled treatment of correlation between Gaia DR2 and DR3 velocity and corrections for the changing light-travel time to this high proper motion system. We revise the planetās mass upward to and find that its orbit is nearly circular and close to coplanar with the outer debris disk. We further present one of the first models of an exoplanet orbit exclusively from absolute astrometry and independently confirm the planetās orbital period. We make specific predictions for the planetās location at key imaging epochs from past and future observing campaigns. We discuss and resolve tensions between previous works regarding the eccentricity, inclination, and mass. Our results further support that Eridani bĀ is one of the closest analogs to a Solar System planet yet detected around a nearby star.
ā±
@@ 1 Introduction
EridaniĀ is one of the closest planetary systems to Earth, just pc away. This proximity has made it a compelling target for direct imaging, giving us an opportunity to potentially image a true analogue to a planet in our Solar System: a relatively mature (200-800 Myr old, Mamajek & Hillenbrand (2008); 439 or 659 Myr old, Barnes (2007)) near-Jupiter mass planet, orbiting a Sun-like star at a separation that would fall between Mars and Jupiter ( AU).
@@
This exosystem has been extensively studied since the 1980s, when pioneering Doppler radial velocity (RV) surveys by Campbell etĀ al. (1988) detected a long-term trend, and then Hatzes etĀ al. (2000) confirmed periodic motion of the star. Further credence was given to these RV signals when Benedict etĀ al. (2006) determined a complete orbit by combining RV data with astrometry from the Hubble Telescopeās Fine Guidance Sensor (FGS), the Allegheny Multichannel Astrometric Photometer (MAP) (Gatewood, 1987, 2000), and Hipparcos (van Leeuwen etĀ al., 1997), making it one of the earliest planets with an orbit derived in-part from absolute astrometry.
@@
Since this time, studies of Eridani bĀ have entered a renaissance, driven by the promise that high-contrast direct imaging technology has advanced to the point where we may, in the near future, be able to study the planet directly via its thermally emitted light.
@@
Several major attempts to image this system in the near-infrared have been published to date, including K-band imaging at Keck by Macintosh etĀ al. (2003), and M-band imaging with MMT by Heinze etĀ al. (2008), L-band imaging with VLT by Janson etĀ al. (2008), L- and M-band imaging with Spitzer by Janson, Markus etĀ al. (2015), M-band imaging with Keck by Mawet etĀ al. (2019), M-band imaging with Keck by Llop-Sayson etĀ al. (2021), and N-band at VLT imaging with Pathak etĀ al. (2021). Though these efforts did not achieve an imaging detection, they placed upper limits of on the brightness of the planet, and therefore limit its mass to for a 400 Myr system age.
@@
Motivated by these attempts, upcoming and completed observations from JWST using NIRCam, MIRI, and NIRSpec are poised to reach significantly better sensitivity. This renews the need to make specific predictions of the planetās position in order to determine the best epochs to observe the planet, measure its mass and thus expected brightness, and confirm the location of any observed point sources. In addition, certain observations could benefit from knowing the location of the planet in advance to adjust the detector placement to mitigate saturation, plan dither patterns, and so on.
@@
The most recent works to re-examine the orbit of Eridani bĀ include Reffert & Quirrenbach (2011), Giguere etĀ al. (2016), Llop-Sayson etĀ al. (2021), Roettenbacher etĀ al. (2022), FritzĀ Benedict (2022), and Feng etĀ al. (2023). These works examined the orbit of Eridani bĀ using a mix of data from various RV sources, Hipparcos (van Leeuwen, 2007), Gaia DR2 (Gaia Collaboration etĀ al., 2018), and Gaia DR3 (Gaia Collaboration etĀ al., 2023), but arrived at somewhat inconsistent conclusions with regard to the planetās orbital eccentricity, inclination, and dynamical mass.
@@
The goal of this paper is to determine the orbit of Eridani bĀ using a careful reanalysis of all available archival data. We combine the superset of all observations considered in previous works, including radial velocity data from eight RV instruments (APF, HIRES, HARPS, Lick, CFHT, CHIRON, EXPRES, and CES) and four astrometric data sources (Hipparcos, Hubble FGS, Gaia DR2, and DR3). Throughout, we take special care to account for systematics and second-order effects that become non-negligible for fast-moving stars over long time-scales, including rigorous propagation of the systemās barycentric motion, perspective acceleration, changing parallax, and changing light travel-time to the receding system. We compare and contrast the methods and results of several previous analyses, finding where they agree and identifying which assumptions had led to disagreements.
@@
In anticipation of a successful imaging detection, Eridani bĀ can also serve as a test run for future deep imaging campaigns with JWST, Roman, or HWO. Some of these campaigns will depend on radial velocity and astrometry to determine the location and brightness of planets ahead of time. Due to the long integrations required to detect sub-Jovian planets, it is incumbent on the community to demonstrate the ability to robustly determine planet positions and brightness ahead of targeted observations. We hope that a correct prediction of the location of Eridani bĀ can demonstrate this capability.
@@ 2.1 Radial Velocity Data
We gathered radial velocity data from many sources, starting with some of the earliest RV measurements made in the 1980s by CFHT (Campbell etĀ al., 1988). We used the ācorrectedā data presented in Table 1. We used data from Lick gathered from Fischer etĀ al. (2014), and split it into four groups as noted by the authors: Lick (1), before November 1994; Lick (2), dewar 13; Lick (3), dewar 6; Lick (4), dewar 8. We used data from CES (Zechmeister etĀ al., 2013), both from the LC and VLC, and treated these data separately. We further included data from HARPS (Trifonov etĀ al., 2020), split between before and the 2015 upgrade (āpreā and āpostā respectively). We used all HIRES and APF RVs presented by Mawet etĀ al. (2019) and by Llop-Sayson etĀ al. (2021). We further added RVs from CHIRON reported by Giguere etĀ al. (2016) and from EXPRES reported by Roettenbacher etĀ al. (2022).
@@
Over such a long time-scale, 3D perspective effects caused by the starās high proper motion ( 1ā/yr) and barycentric radial velocity (16 km/s) have the potential to build up to detectable levels. We elected to handle this apparent acceleration term directly in our models and thus attempted to undo any pre-existing static corrections applied to each dataset. We assumed that this correction was not applied for the CFHT, CES, HIRES, CHIRON, and EXPRES datasets and assumed that the correction was applied for the HIRES, APF, HARPS, and Lick datasets. For the latter datasets, we added back in secular acceleration at the rate given by Gaia DR3. To be conservative, we nonetheless include a linear trend for each dataset in our model. We encourage authors to include descriptions of these corrections in their works so that future authors can make the best use of their data, especially for systems with less orbital coverage than Eridani.
@@
We use the Celerite (Foreman-Mackey etĀ al., 2017) framework through the Celerite.jl
package111https://github.com/ericagol/celerite.jl/ and define an approximate quasi-periodic GP kernel with parameters , , , and , just as in Roettenbacher etĀ al. (2022).
We modeled each dataset with an independent Gaussian process, and multiplied their likelihoods. This treatment inherently assumes that interleaved data from different instruments are uncorrelated (see Blunt etĀ al. (2023), section 4). As discussed in Section 4.3, we tested the impacts of this assumption by comparing against a model with a single GP shared across all instruments, ultimately finding that our results in this case are not sensitive to the exact treatment.
@@ 2.2 Astrometric Data
The EridaniĀ system has been monitored for astrometric motion and parallax for over a century. We include precision absolute astrometry from Hipparcos (van Leeuwen, 2007), Hubble Fine Guidance Sensor (FGS, from Benedict etĀ al., 2006), and both Gaiaās second (DR2, Gaia Collaboration etĀ al., 2018) and third data releases (DR3, Gaia Collaboration etĀ al., 2023).
@@
Beginning with Hipparcos, we used data from the mission in two different ways. We used the Hipparcos IAD epoch astrometry sourced from the Java Tool following the guidance of Nielsen etĀ al. (2020), renormalizing the uncertainties, and testing that we could reproduce the reported catalog measurements and uncertainties exactly. These data were used for the Hipparcos IAD model to be described shortly. We also used the Hipparcos proper motion sourced from the DR3 Hipparcos-Gaia catalog of accelerations (HGCA; Brandt, 2021), calculated from a mix of the 1997 and 2007 Hipparcos data releases, and calibrated against the Gaia DR3 velocity reference frame. These data were used in all other relevant astrometry models to ensure the reference frame was well-calibrated against that of Gaia DR3. This is a key departure from Feng etĀ al. (2023), who did not use the HGCA calibration between Gaia and Hipparcos reference frames, electing to include the reference frame alignment as variables in their model.
@@
The second source of absolute astrometry data we used for this system was that of Benedict etĀ al. (2006). The data used in that publication were derived by measuring the positions of EridaniĀ using the Hubble Fine Guidance Sensor (FGS) relative to a field of nearby background stars. Although the raw data are available, the re-reduction of these data is out of scope for this work. Instead, we opted to extract the residuals from their parallax and proper motion fit (Figure 7 of Benedict etĀ al. (2006)) using Web Plot Digitizer (Rohatgi, 2025).
@@
The extracted residuals are relative to a five-parameter linear model of annual parallax and proper motion fit to a combination of the FGS data, lower-precision astrometry from the Allegheny Multichannel Astrometric Photometer (not considered in this work). We justify using these residuals in our model by noting that the authorsā best-fit five-parameter model arrived at a parallax value that is very close to the measurement from Gaia DR3, meaning that no annual cyclic motion should be detectable in the residuals. Then, any residual proper motion not captured by their best-fit five-parameter model would appear as a linear trend. Such a linear trend could also be caused by frame rotation versus the Gaia DR3 reference frame, which we must assume is present to some degree. Rather than attempt to calibrate the absolute astrometry against Gaia with sufficient accuracy, we elect to simply fit linear trends to the residuals and fit only the acceleration over the FGS observations, which will be the same in any inertial reference frame. This is analogous to including a linear trend in models of RV data. This neatly sidesteps the issue of determining an absolute calibration of the FGS data and allowing us to use the residuals directly.
@@
We model the two catalogs by simulating Gaia observations at each visibility window reported by GOST 222https://gaia.esac.esa.int/gost/index.jsp. We approximate the data using a single measurement per visibility window rather than per scan, just as in most previous works. We then perform a linear least-squares fit of a 4-parameter model to both the DR2 time-range and the DR3 time-range. We then compared the velocities extracted from these simulations, against the reported catalog values .
This is similar in principle to the work of (Feng etĀ al., 2023), though we differ in how the data are modeled a few important respects. First, we handle the correlation between DR2 and DR3, and second, we use the HGCAās spatially dependent calibration between the Hipparcos and Gaia reference frames and calibrated (inflated) uncertainties rather than those reported by Gaia directly.
@@
One additional consideration is that the DR2 proper motion reference frame is known to be rotating slightly compared to DR3. We corrected for this according to from Lindegren, Lennart (2020). Unlike previous works that simply adopted the mean values of these parameters, we included them as priors in our models in order to rigorously marginalize over the uncertainties in the reference frame calibration. This is similar in effect to the offsets employed by Feng etĀ al. (2023), but only applied to the DR2-DR3 velocity difference.
As noted by the Gaia teams and in the HGCA (Gaia Collaboration etĀ al., 2016; Brandt, 2018, 2021), the formal uncertainties reported by Gaia and Hipparcos are underestimated. Accordingly, the the HGCA catalogs inflate the proper motion uncertainties such that a sample stars that are not likely to be accelerating have standardized residuals that follow a Gaussian distribution. This error inflation factor for Gaia DR2 and DR3 varies with spatial position. Where Feng etĀ al. (2023) adopts a flexible jitter parameter, we used the calibrated uncertainties from the HGCA catalogs for both DR2 and DR3.
@@ 2.3 Imaging
There is raw image data available from previous campaigns, including those by Mawet etĀ al. (2019), Llop-Sayson etĀ al. (2021), and Pathak etĀ al. (2021). We did not include these images in our model since the published sensitivity curves indicate that the planet would be below the level of the noise for reasonable inclination angles (limit of from Mawet etĀ al. (2019), 2ā4 Ā from Pathak etĀ al. (2021)). A future work could explore integrating image data into this orbit model following the same recipes as Llop-Sayson etĀ al. (2021) and Thompson etĀ al. (2022), especially if more images of comparable or greater sensitivity become available.

### v2

@@ Abstract
The mature Jovian planet Eridani b orbits one of the closest sun-like stars at a moderate separation of 3.5 AU, presenting one of the best opportunities to image a true analog to a solar system planet. We perform a thorough joint reanalysis and cross-validation of all available archival radial velocity and astrometry data, combining data from eight radial velocity instruments and four astrometric sources (Hipparcos, Hubble FGS, Gaia DR2, and Gaia DR3). We incorporate methodological advances that impact our findings including a principled treatment of correlation between Gaia DR2 and DR3 velocity. We revise the planet’s mass upward to and find that its orbit is nearly circular and likely to be to coplanar with the outer debris disk. We further present one of the first models of an exoplanet orbit exclusively from absolute astrometry and independently confirm the planet’s orbital period. We make specific predictions for the planet’s location at key imaging epochs from past and future observing campaigns. We discuss and resolve tensions between previous works regarding the eccentricity, inclination, and mass. Our results further support that Eridani b is one of the closest analogs to a Solar System planet yet detected around a nearby star.
@@ 1 Introduction
Eridani is one of the closest planetary systems to Earth, just pc away. This proximity has made it a compelling target for direct imaging, giving us an opportunity to potentially image a true analogue to a planet in our Solar System: a relatively mature (200-800 Myr old, Mamajek & Hillenbrand (2008); 439 or 659 Myr old, Barnes (2007)) near-Jupiter mass planet, orbiting a Sun-like star at a separation that would fall between Mars and Jupiter ( AU).
@@
This exosystem has been extensively studied since the 1980s, when pioneering Doppler radial velocity (RV) surveys by Campbell et al. (1988) detected a long-term trend, and then Hatzes et al. (2000) confirmed periodic motion of the star. Further credence was given to these RV signals when Benedict et al. (2006) determined a complete orbit by combining RV data with astrometry from the Hubble Telescope’s Fine Guidance Sensor (FGS), the Allegheny Multichannel Astrometric Photometer (MAP) (Gatewood, 1987, 2000), and Hipparcos (van Leeuwen et al., 1997), making it one of the earliest planets with an orbit derived in-part from absolute astrometry.
@@
Since this time, studies of Eridani b have entered a renaissance, driven by the promise that high-contrast direct imaging technology has advanced to the point where we may, in the near future, be able to study the planet directly via its thermally emitted light.
@@
Several major attempts to image this system in the near-infrared have been published to date, including K-band imaging at Keck by Macintosh et al. (2003), and M-band imaging with MMT by Heinze et al. (2008), L-band imaging with VLT by Janson et al. (2008), L- and M-band imaging with Spitzer by Janson, Markus et al. (2015), M-band imaging with Keck by Mawet et al. (2019), M-band imaging with Keck by Llop-Sayson et al. (2021), and N-band at VLT imaging with Pathak et al. (2021). Though these efforts did not achieve an imaging detection, they placed upper limits of on the brightness of the planet, and therefore limit its mass to for a 400 Myr system age.
@@
Motivated by these attempts, upcoming and completed observations from JWST using NIRCam, MIRI, and NIRSpec are poised to reach significantly better sensitivity. This renews the need to make specific predictions of the planet’s position in order to determine the best epochs to observe the planet, measure its mass and thus expected brightness, and confirm the location of any observed point sources. In addition, certain observations could benefit from knowing the location of the planet in advance to adjust the detector placement to mitigate saturation, plan dither patterns, and so on.
@@
The most recent works to re-examine the orbit of Eridani b include Reffert & Quirrenbach (2011), Giguere et al. (2016), Llop-Sayson et al. (2021), Roettenbacher et al. (2022), Fritz Benedict (2022), and Feng et al. (2023). These works examined the orbit of Eridani b using a mix of data from various RV sources, Hipparcos (van Leeuwen, 2007), Gaia DR2 (Gaia Collaboration et al., 2018), and Gaia DR3 (Gaia Collaboration et al., 2023), but arrived at somewhat inconsistent conclusions with regard to the planet’s orbital eccentricity, inclination, and dynamical mass.
@@
The goal of this paper is to determine the orbit of Eridani b using a careful reanalysis of all available archival data. We combine the superset of all observations considered in previous works, including radial velocity data from eight RV instruments (APF, HIRES, HARPS, Lick, CFHT, CHIRON, EXPRES, and CES) and four astrometric data sources (Hipparcos, Hubble FGS, Gaia DR2, and DR3). Throughout, we take special care to account for systematics and second-order effects that become non-negligible for fast-moving stars over long time-scales, including rigorous propagation of the system’s barycentric motion, perspective acceleration, changing parallax, and changing light travel-time to the receding system. We compare and contrast the methods and results of several previous analyses, finding where they agree and identifying which assumptions had led to disagreements.
@@
In anticipation of a successful imaging detection, Eridani b can also serve as a test run for future deep imaging campaigns with JWST, Roman, or HWO. Some of these campaigns will depend on radial velocity and astrometry to determine the location and brightness of planets ahead of time. Due to the long integrations required to detect sub-Jovian planets, it is incumbent on the community to demonstrate the ability to robustly determine planet positions and brightness ahead of targeted observations. We hope that a correct prediction of the location of Eridani b can demonstrate this capability.
@@ 2.1 Radial Velocity Data
We gathered radial velocity data from many sources, starting with some of the earliest RV measurements made in the 1980s by CFHT (Campbell et al., 1988). We used the “corrected” data presented in Table 1. We used data from Lick gathered from Fischer et al. (2014), and split it into four groups as noted by the authors: Lick (1), before November 1994; Lick (2), dewar 13; Lick (3), dewar 6; Lick (4), dewar 8. We used data from CES (Zechmeister et al., 2013), both from the LC and VLC, and treated these data separately. We further included data from HARPS (Trifonov et al., 2020), split between before and the 2015 upgrade (“pre” and “post” respectively). We used all HIRES and APF RVs presented by Mawet et al. (2019) and by Llop-Sayson et al. (2021). We further added RVs from CHIRON reported by Giguere et al. (2016) and from EXPRES reported by Roettenbacher et al. (2022).
@@
Over such a long time-scale, 3D perspective effects caused by the star’s high proper motion ( 1”/yr) and barycentric radial velocity (16 km/s) have the potential to build up to detectable levels. We elected to handle this apparent acceleration term directly in our models and thus attempted to undo any pre-existing static corrections applied to each dataset. We assumed that this correction was not applied for the CFHT, CES, HIRES, CHIRON, and EXPRES datasets and assumed that the correction was applied for the HIRES, APF, HARPS, and Lick datasets. For the latter datasets, we added back in secular acceleration at the rate given by Gaia DR3. To be conservative, we nonetheless include a linear trend for each dataset in our model. We encourage authors to include descriptions of these corrections in their works so that future authors can make the best use of their data, especially for systems with less orbital coverage than Eridani.
@@
We use the Celerite (Foreman-Mackey et al., 2017) framework through the Celerite.jl package111https://github.com/ericagol/celerite.jl/ and define an approximate quasi-periodic GP kernel with parameters , , , and , just as in Roettenbacher et al. (2022).
We modeled each dataset with an independent Gaussian process, and multiplied their likelihoods. This treatment inherently assumes that interleaved data from different instruments are uncorrelated (see Blunt et al. (2023), section 4). As discussed in Section 4.3, we tested the impacts of this assumption by comparing against a model with a single GP shared across all instruments, ultimately finding that our results in this case are not sensitive to the exact treatment.
@@ 2.2 Astrometric Data
The Eridani system has been monitored for astrometric motion and parallax for over a century. We include precision absolute astrometry from Hipparcos (van Leeuwen, 2007), Hubble Fine Guidance Sensor (FGS, from Benedict et al., 2006), and both Gaia’s second (DR2, Gaia Collaboration et al., 2018) and third data releases (DR3, Gaia Collaboration et al., 2023).
@@
Beginning with Hipparcos, we used data from the mission in two different ways. We used the Hipparcos IAD epoch astrometry sourced from the Java Tool following the guidance of Nielsen et al. (2020), renormalizing the uncertainties, and testing that we could reproduce the reported catalog measurements and uncertainties exactly. These data were used for the Hipparcos IAD model to be described shortly. We also used the Hipparcos proper motion sourced from the DR3 Hipparcos-Gaia catalog of accelerations (HGCA; Brandt, 2021), calculated from a mix of the 1997 and 2007 Hipparcos data releases, and calibrated against the Gaia DR3 velocity reference frame. These data were used in all other relevant astrometry models to ensure the reference frame was well-calibrated against that of Gaia DR3. This is a key departure from Feng et al. (2023), who did not use the HGCA calibration between Gaia and Hipparcos reference frames, electing to include the reference frame alignment as variables in their model.
@@
The second source of absolute astrometry data we used for this system was that of Benedict et al. (2006). The data used in that publication were derived by measuring the positions of Eridani using the Hubble Fine Guidance Sensor (FGS) relative to a field of nearby background stars. Although the raw data are available, the re-reduction of these data is out of scope for this work. Instead, we opted to extract the residuals from their parallax and proper motion fit (Figure 7 of Benedict et al. (2006)) using Web Plot Digitizer (Rohatgi, 2025).
@@
The extracted residuals are relative to a five-parameter linear model of annual parallax and proper motion fit to a combination of the FGS data, lower-precision astrometry from the Allegheny Multichannel Astrometric Photometer (not considered in this work). We justify using these residuals in our model by noting that the authors’ best-fit five-parameter model arrived at a parallax value that is very close to the measurement from Gaia DR3, meaning that no annual cyclic motion should be detectable in the residuals. Then, any residual proper motion not captured by their best-fit five-parameter model would appear as a linear trend. Such a linear trend could also be caused by frame rotation versus the Gaia DR3 reference frame, which we must assume is present to some degree. Rather than attempt to calibrate the absolute astrometry against Gaia with sufficient accuracy, we elect to simply fit linear trends to the residuals and fit only the acceleration over the FGS observations, which will be the same in any inertial reference frame. This is analogous to including a linear trend in models of RV data. This neatly sidesteps the issue of determining an absolute calibration of the FGS data and allowing us to use the residuals directly.
@@
We model the two catalogs by simulating Gaia observations at each visibility window reported by GOST 222https://gaia.esac.esa.int/gost/index.jsp. We remove known gaps in Gaia data, leaving only one extra epoch from GOST unaccounted for. We include this visibility window as a discrete parameter in our model and marginalize over it.
We then perform a linear least-squares fit of a 5-parameter model to both the DR2 time-range and the DR3 time-range. We then compared the velocities extracted from these simulations, against the reported catalog values .
This is similar in principle to the work of (Feng et al., 2023), though we differ in how the data are modeled a few important respects. First, we handle the correlation between DR2 and DR3, and second, we use the HGCA’s spatially dependent calibration between the Hipparcos and Gaia reference frames and calibrated (inflated) uncertainties rather than those reported by Gaia directly combined with flexible jitter parameters.
@@
One additional consideration is that the DR2 proper motion reference frame is known to be rotating slightly compared to DR3. We corrected for this according to from Lindegren, Lennart (2020). Unlike previous works that simply adopted the mean values of these parameters, we included them as priors in our models in order to rigorously marginalize over the uncertainties in the reference frame calibration. This is similar in effect to the offsets employed by Feng et al. (2023), but only applied to the DR2-DR3 velocity difference.
As noted by the Gaia teams and in the HGCA (Gaia Collaboration et al., 2016; Brandt, 2018, 2021), the formal uncertainties reported by Gaia and Hipparcos are underestimated. Accordingly, the the HGCA catalogs inflate the proper motion uncertainties such that a sample stars that are not likely to be accelerating have standardized residuals that follow a Gaussian distribution. This error inflation factor for Gaia DR2 and DR3 varies with spatial position. Where Feng et al. (2023) adopts a flexible jitter parameter, we used the calibrated uncertainties from the HGCA catalogs for both DR2 and DR3.
@@ 2.3 Imaging
There is raw image data available from previous campaigns, including those by Mawet et al. (2019), Llop-Sayson et al. (2021), and Pathak et al. (2021). We did not include these images in our model since the published sensitivity curves indicate that the planet would be below the level of the noise for reasonable inclination angles (limit of from Mawet et al. (2019), 2–4 from Pathak et al. (2021)). A future work could explore integrating image data into this orbit model following the same recipes as Llop-Sayson et al. (2021) and Thompson et al. (2022), especially if more images of comparable or greater sensitivity become available.